the heat of vaporization of water is 40.66 kJ/mol. How much heat is absorbed when 2.88 g of water boils at atmospheric pressure? ____kJ

(show wrk)
Answer #1
Concepts and reason

The relationship between specific heat of a substance and heat gained or lost is as follows:

Q=n×ΔHvap{\rm{Q = n}}{\kern 1pt} {\kern 1pt} {\rm{ \times }}{\kern 1pt} {\kern 1pt} {\kern 1pt} \Delta {{\rm{H}}_{{\rm{vap}}}}

Here, n is number of moles of the substance.

ΔHvap\Delta {{\rm{H}}_{{\rm{vap}}}} is the heat vaporization of the substance.

Fundamentals

Heat vaporization is of a substance is the amount of heat required to vaporize one mole of the substance.

Calculate the number of moles of water as follows:

numberofmolesofwater=massmolarmass=2.88g18.0g/mol=0.16mol\begin{array}{l}\\{\rm{number}}{\kern 1pt} {\kern 1pt} {\rm{of}}{\kern 1pt} {\kern 1pt} {\rm{moles}}{\kern 1pt} {\kern 1pt} {\kern 1pt} {\rm{of}}{\kern 1pt} {\kern 1pt} {\rm{water}}{\kern 1pt} {\kern 1pt} {\kern 1pt} {\rm{ = }}{\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} \frac{{{\rm{mass}}}}{{{\rm{molar}}{\kern 1pt} {\kern 1pt} {\kern 1pt} {\rm{mass}}}}\\\\{\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} = {\kern 1pt} {\kern 1pt} {\kern 1pt} \frac{{2.88{\kern 1pt} {\kern 1pt} {\rm{g}}}}{{18.0{\kern 1pt} {\kern 1pt} {\rm{g/mol}}}}\\\\{\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} = {\kern 1pt} {\kern 1pt} 0.16{\kern 1pt} {\kern 1pt} {\rm{mol}}\\\end{array}

Now, calculate the heat energy absorbed to boil the water at atmospheric pressure is as follows:

Q=n×ΔHvap=0.16mol×40.66kJ/mol=6.506kJ\begin{array}{l}\\{\rm{Q = n}}{\kern 1pt} {\kern 1pt} {\rm{ \times }}{\kern 1pt} {\kern 1pt} {\kern 1pt} \Delta {{\rm{H}}_{{\rm{vap}}}}{\rm{ }}\\\\{\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\rm{ = }}{\kern 1pt} {\kern 1pt} 0.16{\kern 1pt} {\kern 1pt} {\rm{mol}} \times 40.66{\kern 1pt} {\kern 1pt} {\rm{kJ/mol}}\\\\{\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\kern 1pt} {\rm{ = }}{\kern 1pt} {\kern 1pt} 6.506{\kern 1pt} {\kern 1pt} {\rm{kJ}}\\\end{array}

Therefore, the total heat absorbed is 6.506kJ6.506{\kern 1pt} {\kern 1pt} {\rm{kJ}} .

Ans:

The heat energy absorbed is 6.506kJ6.506{\kern 1pt} {\kern 1pt} {\rm{kJ}} .